Measures of liquidity, Solvency, and Profitability

The comparative financial statements of Marshall Inc. are as follows. The market price of Marshall common stock was $ 58 on December 31, 20Y2.

Marshall Inc.
Comparative Retained Earnings Statement
For the Years Ended December 31, 20Y2 and 20Y1
   20Y2    20Y1
Retained earnings, January 1 $1,802,125 $1,537,175
Net income 425,600 314,800
Total $2,227,725 $1,851,975
Dividends:
On preferred stock $13,300 $13,300
On common stock 36,550 36,550
Total dividends $49,850 $49,850
Retained earnings, December 31 $2,177,875 $1,802,125
Marshall Inc.
Comparative Income Statement
For the Years Ended December 31, 20Y2 and 20Y1
   20Y2    20Y1
Sales $2,628,000 $2,421,260
Cost of goods sold 927,830 853,600
Gross profit $1,700,170 $1,567,660
Selling expenses $587,930 $723,500
Administrative expenses 500,820 424,920
Total operating expenses $1,088,750 $1,148,420
Income from operations $611,420 $419,240
Other revenue 32,180 26,760
$643,600 $446,000
Other expense (interest) 160,000 88,000
Income before income tax $483,600 $358,000
Income tax expense 58,000 43,200
Net income $425,600 $314,800
Marshall Inc.
Comparative Balance Sheet
December 31, 20Y2 and 20Y1
   20Y2    20Y1
Assets
Current assets
Cash $549,670 $432,510
Marketable securities 831,930 716,720
Accounts receivable (net) 452,600 423,400
Inventories 335,800 262,800
Prepaid expenses 103,990 86,500
Total current assets $2,273,990 $1,921,930
Long-term investments 802,973 425,709
Property, plant, and equipment (net) 2,600,000 2,340,000
Total assets $5,676,963 $4,687,639
Liabilities
Current liabilities $689,088 $975,514
Long-term liabilities:
Mortgage note payable, 8% $900,000 $0
Bonds payable, 8% 1,100,000 1,100,000
Total long-term liabilities $2,000,000 $1,100,000
Total liabilities $2,689,088 $2,075,514
Stockholders' Equity
Preferred $0.70 stock, $20 par $380,000 $380,000
Common stock, $10 par 430,000 430,000
Retained earnings 2,177,875 1,802,125
Total stockholders' equity $2,987,875 $2,612,125
Total liabilities and stockholders' equity $5,676,963 $4,687,639

Required:

Determine the following measures for 20Y2, rounding to one decimal place, except for dollar amounts, which should be rounded to the nearest cent. Use the rounded answer of the requirement for subsequent requirement, if required. Assume 365 days a year.

10. Times interest earned
11. Asset turnover
12. Return on total assets %
13. Return on stockholders’ equity %
14. Return on common stockholders’ equity %
15. Earnings per share on common stock $
16. Price-earnings ratio
17. Dividends per share of common stock $
18. Dividend yield %
Answer #1

Sulotion:

10) Times interest earned

= EBIT ÷ Interest Expense

= 611420 ÷ 160000

= 3.82 times

11) Assets Turn over ratio

=Sales ÷ Total assets

= 2628000 ÷ 5676963

= 0.462923 times

12) Return on total assets

= (EBIT or Income from operation ÷ Total assets)× 100

= ( 611420 ÷ 5676963) × 100

= 10.77 % approximately

13) Return on stock holders equity

=( Net Income ÷ Total stock holders equity) × 100

=( 425600 ÷ 2987875) × 100

= 14.24% approximately